A radioactive sample of 238U decays to Pb through a process for which the half-life is 4.5 x 109 years. Find the ratio of number of nuclei of Pb to 238U after a time of 1.5 x 109 years. Given (2)1/3 = 1.26 .

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Hello Student,
Please find the answer to your question
KEY CONCEPT:
The wavelength λ, of photon for different lines of Balmer series is given by
hc / λ = 13.6 [ 1 / 22 – 1 / n2] eV, where n = 3, 4, 5
Using above relation, we get the value of λ = 657 nm, 487 nm between 450 nm and 700 nm. Since 487 nm, is smaller than 657 nm electron of max. E.K. will be emitted for photon corresponding to wavelength 487 nm with
(K.E.) = hc / λ – W = ( 1242 / 487 - 2) = 0.55 eV
Thanks
Navjot Kalra
askIITians Faculty